# Numerical valuation of European options under two-asset infinite-activity exponential Lévy models

### v1

@@ 3.2.1 Diffusion term
$\beta_{m}^{\left(-1\right)}=\frac{2}{h_{x,m}\left(h_{x,m}+h_{x,m+1}\right)},\qquad\beta_{m}^{\left(0\right)}=\frac{-2}{h_{x,m}h_{x,m+1}},\qquad\beta_{m}^{\left(1\right)}=\frac{2}{h_{x,m+1}\left(h_{x,m}+h_{x,m+1}\right)}.$
Concerning the boundary of the truncated spatial domain, we modify the previous formulae in the following way. At the lower boundary $x_{0}=0$, the first- and second-order derivative terms in (3.8) vanish. Hence, it is natural to choose $\alpha_{0}^{\left(j\right)}=0$ and $\beta_{0}^{\left(j\right)}=0$ for any $j=\left\{-1,0,1\right\}$. At the upper boundary $x_{N_{x}}=x_{\max}$, we make the natural assumption that the solution $v$ behaves linearly in $x$, thus we choose $\beta_{N_{x}}^{\left(j\right)}=0$ for any $j=\left\{-1,0,1\right\}$, and we approximate the first-order derivative by the first-order backward finite difference scheme.
@@ 3.2.2 Summation term
    Here, $0_{P\times Q}$ denotes the null matrix of dimensions $P\times Q$, $\text{vec}\left(\cdot\right)$ denotes the vectorization of a matrix, $\sharp$ indicates the number of points of a given grid in one direction and $\Omega\in\mathbb{R}^{\sharp\mathbf{z}\times\sharp\mathbf{z}}$ is the matrix whose entries are the coefficients $\omega_{l_{1}l_{2}}$ defined by (3.5). For an example of a matrix $C$, we refer to Appendix B.
@@ 3.3 Temporal discretization: the semi-Lagrangian $\theta$-method
$Y^{n,0}=\begin{cases}V^{n-1}&n=1,\\ 2V^{n-1}-V^{n-2}&n=2,\\ 3V^{n-1}-3V^{n-2}+V^{n-3}&n=3,\\ 4V^{n-1}-6V^{n-2}+4V^{n-3}-V^{n-4}&n\geq 4.\end{cases}$ (3.29)
@@ 4 Numerical experiments
$\varphi\left(\xi\right)=\begin{cases}c\xi&0\leq\xi\leq\xi_{\rm int},\\ x_{\rm int}+c\sinh\left(\xi-\xi_{\rm int}\right)&\xi_{\rm int}<\xi\leq\xi_{\rm max},\end{cases}$
@@
where $K_{\nu}\left(\tau\right)=\frac{1}{2}\int_{0}^{\infty}y^{\nu-1}e^{-\frac{1}{2}\tau\left(y+y^{-1}\right)}dy$, for $\tau>0$, denotes the modified Bessel function of the second kind, [^6] See Schoutens (2003, Appendix A). $\left\langle x,y\right\rangle_{\rho}=x^{\top}\rho^{-1}y$ and $\left\|x\right\|_{\rho}=\sqrt{\left\langle x,x\right\rangle_{\rho}}$ is its induced norm. The constants $A_{\ell},B_{\ell}$ and $C_{\ell}$ in (2.5) are defined, with respect to $\left\|\cdot\right\|_{\rho}$, as
$\left(A_{\ell},B_{\ell},C_{\ell}\left(h\right)\right)=\left(2\alpha,\sqrt{\left\|\eta\right\|_{\rho}^{2}+2\lambda}-\left\|\eta\right\|_{\rho},\frac{2^{\alpha}\delta\Gamma\left(1+\alpha\right)}{\pi\sqrt{\det\left[\rho\right]}}e^{h\left\|\eta\right\|_{\rho}}\right).$
@@
    $z_{\max}^{\mathbf{II}}=\sqrt{0.1}z_{\max}^{\mathbf{III}}$. In this way, the size of $R_{z}^{\mathbf{II}}$ is about 10% of the full integration domain $R_{z}$.
@@
    $z_{\max}^{\mathbf{III}}=\max\left\{\left\|z\right\|_{\infty}:z\in\mathbb{R}^{2},\ell\left(z\right)=10^{-8}\right\}$. Since the Lévy measure decays at least exponentially as $\left\|z\right\|\rightarrow\infty$, we ensure that $\ell\left(z\right)<10^{-8}$ for all $z\in\mathbb{R}^{2}$ such that $\left\|z\right\|_{\infty}>z_{\max}^{\mathbf{III}}$.
@@
    $c$ is chosen such that $F=\max\left(65\%,\frac{x_{\rm int}}{x_{\rm max}}\right)$. In this way, approximately at least 65% of the spatial grid points in each given direction are in the interval $\left[0,x_{\rm int}\right]$.
@@
    $N_{y}^{-}=\text{ceil}\left[-\frac{1}{h_{z}}\ln\left(x_{1}\right)\right]+N_{y}^{*}$ and $N_{y}^{+}=\text{ceil}\left[\frac{1}{h_{z}}\ln\left(x_{\rm max}\right)\right]+N_{y}^{*}$ for some given $N_{y}^{*}\in\mathbb{N}_{0}$. This choice minimises the need for extrapolation in (3.17) as it is necessary to extrapolate just to the grid points $x_{m_{1}m_{2}}$ with either $m_{1}=0$ or $m_{2}=0$. This is done in a linear fashion. In (3.16), we set $v\left(x,t\right)=0$ whenever $x\notin R_{x}$.
@@
    $N_{y}^{*}$ is taken as the minimal $n\in\mathbb{N}_{0}$ such that the maximal prime factor of $\sharp\textrm{in}=N_{y}^{-}+N_{y}^{+}+2N_{z}$ is at most 7. This is beneficial for the efficiency of the FFT.
@@
    The tolerances used for the fixed-point iteration and BiCGSTAB are set to $10^{-7}$ and $10^{-14}$, respectively.
@@
$\left(90,90\right)$ 12.6534 10.1080 11.4067 11.5830
$\left(90,100\right)$ 10.6121 5.8462 7.8724 8.1529
$\left(90,110\right)$ 9.0136 3.0178 5.1023 5.4657
$\left(100,90\right)$ 10.4061 5.7637 7.8897 8.0910
$\left(100,100\right)$ 8.8015 2.9037 5.1186 5.3953
$\left(100,110\right)$ 7.5309 1.3893 3.1156 3.4311
$\left(110,90\right)$ 8.6181 2.8070 5.1393 5.3381
$\left(110,100\right)$ 7.3464 1.3062 3.1326 3.3735
$\left(110,110\right)$ 6.3290 0.6014 1.7937 2.0397
provides the numerical option prices for various points $x$ around $\left(K,K\right)$ and all four parameter sets from Table 1.
We next investigate the convergence behaviour of the numerical scheme. Let $\mathbf{x}_{N}$ denote the set of spatial grid points if $N_{x}=N$. For $x\in R_{x}$, let $\tilde{u}(x;N)$ denote the approximation of the exact solution value $u(x,T)$ obtained by the numerical scheme if $N_{x}=N$. More precisely, the vector $V^{N_{t}}$ generated by (3.27), (3.28) yields the approximation on the spatial grid $\mathbf{x}_{N}$ and cubic interpolation is employed whenever $x\notin\mathbf{x}_{N}$. We consider $\tilde{u}(x;N)$ with $N=400$ as the reference solution and study for $50\leq N\leq 200$ the total error defined by
@@
displays the total errors for all four parameter sets from Table 1. The quantity $p$ in the legend denotes the numerical order of convergence, which is computed by linear regression. Clearly, the favourable result is found that the numerical scheme achieves second-order convergence for each set of parameters.
@@ 5 Conclusions
In this paper, we have developed an effective numerical method for the valuation of European options under two-asset exponential Lévy models with particular attention to the infinite-activity case. Our method is based upon the ideas in Wang et al. (2007) for the one-asset case. A key part of our method is the tailored discretization of the non-local integral term, designed to handle singular measures under mild assumptions. The discretized integral term can subsequently be efficiently evaluated by FFT. For the discretization in time, the semi-Lagrangian Crank–Nicolson method is employed with a fixed-point iteration on the integral part. Numerical experiments for put-on-the-average options under Normal Tempered Stable processes indicate that our method achieves favourable second-order convergence whenever the exponential Lévy model has finite-variation.
@@ Appendix A $d$-dimensional Normal Tempered Stable process
In our context, we will use this process to define the jump component of the logarithmic return in asset prices, i.e. we choose
@@ A.1 Tempered Stable subordinator
Lévy measure $\ell_{G}\left(x\right)=\mathbb{I}_{x>0}\delta e^{-\lambda x}x^{-1-\alpha}$
Characteristic exponent $\psi_{G}\left(x\right)=\begin{cases}-\delta\ln\left(1-ix\lambda^{-1}\right)&\text{if }\alpha=0\\ \delta\Gamma\left(-\alpha\right)\left(\left(\lambda-ix\right)^{\alpha}-\lambda^{\alpha}\right)&\text{if }\alpha\in\left(0,1\right)\end{cases}$
@@
Density function $f_{G}\left(x\right)=\begin{cases}\mathbb{I}_{x>0}\frac{\lambda^{\delta}}{\Gamma\left(\delta\right)}x^{\delta-1}e^{-\lambda x}&\text{if }\alpha=0\\ \mathbb{I}_{x>0}\delta x^{-\frac{3}{2}}e^{-\left(\sqrt{\lambda}x-\delta\sqrt{\pi}\right)^{2}x^{-1}}&\text{if }\alpha=\frac{1}{2}\\ \text{not known analytically}&\text{else}\end{cases}$
@@ A.2 Normal Tempered Stable process
Consider $B\left(t\right)=\eta t+\sqrt{\rho}\,W\left(t\right)$, where $W$ is a standard $d$-dimensional Wiener process, $\eta\in\mathbb{R}^{d}$ and $\sqrt{\rho}$ is the Cholesky decomposition of a given positive semi-definite symmetric matrix $\rho$, i.e., $\rho=\sqrt{\rho}\cdot\sqrt{\rho}^{\top}$. Adapting the results presented in Barndorff-Nielsen et al. (2001) and Rocha-Arteaga & Sato (2019, Chapter 4) [^8] The authors consider the more general case where the characteristic exponent of $L$ is defined as $\psi_{L}\left(\tau\right)=\int_{\mathbb{R}^{d}}\left(e^{i\tau^{\top}z}-1-i\tau^{\top}z\mathbb{I}_{\left\|z\right\|<1}\right)\ell_{L}\left(dz\right)$, while we consider the case where $\psi_{L}\left(\tau\right)=\int_{\mathbb{R}^{d}}\left(e^{i\tau^{\top}z}-1-i\tau^{\top}z\right)\ell_{L}\left(dz\right)$., we define a $d$-dimensional Normal Tempered Stable process as
@@
Lévy measure $\ell_{L}\left(x\right)=\delta\Phi\left(x\mid\alpha,0\right)$
Characteristic exponent $\psi_{L}\left(x\right)=\begin{cases}-\delta\ln\left(\frac{\lambda-ix^{\top}\eta+\frac{1}{2}x^{\top}\rho x}{\lambda}\right)-ix^{\top}c&\text{if }\alpha=0\\ \delta\Gamma\left(-\alpha\right)\left(\left(\lambda-ix^{\top}\eta+\frac{1}{2}x^{\top}\rho x\right)^{\alpha}-\lambda^{\alpha}\right)-ix^{\top}c&\text{if }\alpha\in\left(0,1\right)\end{cases}$
@@
Density function $f_{L}\left(x\right)=\begin{cases}\frac{\lambda^{\delta}}{\Gamma\left(\delta\right)}\Phi\left(x+c\mid-\delta,0\right)&\text{if }\alpha=0\\ \delta e^{2\delta\sqrt{\lambda\pi}}\Phi\left(x+c\mid\frac{1}{2},\delta^{2}\pi\right)&\text{if }\alpha=\frac{1}{2}\\ \text{not known analytically}&\text{else}\end{cases}$
@@
$\Phi\left(x\mid a,b\right)=2\sqrt{\frac{\left(\left\|\eta\right\|_{\rho}^{2}+2\lambda\right)^{a+\frac{d}{2}}}{\left(2\pi\right)^{d}\det\left[\rho\right]}}\frac{K_{a+\frac{d}{2}}\left(\sqrt{\left(\left\|\eta\right\|_{\rho}^{2}+2\lambda\right)\left(\left\|x\right\|_{\rho}^{2}+2b\right)}\right)}{\left(\sqrt{\left\|x\right\|_{\rho}^{2}+2b}\right)^{a+\frac{d}{2}}}e^{\left\langle\eta,x\right\rangle_{\rho}}$ (A.2)
where $K_{\nu}\left(\tau\right)=\frac{1}{2}\int_{0}^{\infty}y^{\nu-1}e^{-\frac{1}{2}\tau\left(y+y^{-1}\right)}dy$, for $\tau>0$, denotes the modified Bessel function of the second kind (see Schoutens (2003, Appendix A)), $\left\langle x,y\right\rangle_{\rho}=x^{\top}\rho^{-1}y$ and $\left\|x\right\|_{\rho}=\sqrt{\left\langle x,x\right\rangle}$ is its induced norm. We conclude this appendix with the following proposition.
@@
Consider a Lévy measure $\ell$ over $\mathbb{R}^{d}_{*}=\mathbb{R}^{d}\setminus\{0\}$. Assume that there exist constants $A_{\ell}$ and $B_{\ell}$, and for any given $h>0$ a constant $C_{\ell}(h)$ such that
$\begin{cases}\ell\left(z\right)\leq C_{\ell}\left(h\right)\left\|z\right\|_{\rho}^{-A_{\ell}-d}&\text{for any }z\in\mathbb{R}^{d}_{*}\text{ such that }\left\|z\right\|_{\rho}\in\left(0,h\right]\\ \ell\left(z\right)=O\left(e^{-B_{\ell}\left\|z\right\|_{\rho}}\right)&\text{as }\left\|z\right\|_{\rho}\rightarrow\infty.\end{cases}$
@@
$A_{\ell}$ $=2\alpha$
$B_{\ell}$ $=\sqrt{\left\|\eta\right\|_{\rho}^{2}+2\lambda}-\left\|\eta\right\|_{\rho}$
$C_{\ell}\left(h\right)$ $=\frac{2^{\alpha}\delta\Gamma\left(\alpha+\frac{d}{2}\right)}{\sqrt{\pi^{d}\det\left[\rho\right]}}e^{h\left\|\eta\right\|_{\rho}}.$

### v2

@@ 3.2.1 Diffusion term
$\beta_{m}^{\left(-1\right)}=\frac{2}{h_{x,m}\left(h_{x,m}+h_{x,m+1}\right)},\qquad\beta_{m}^{\left(0\right)}=\frac{-2}{h_{x,m}h_{x,m+1}},\qquad\beta_{m}^{\left(1\right)}=\frac{2}{h_{x,m+1}\left(h_{x,m}+h_{x,m+1}\right)},$
where $h_{x,m}=x_{m}-x_{m-1}$. Concerning the boundary of the truncated spatial domain, we modify the previous formulae in the following way. At the lower boundary $x_{0}=0$, the first- and second-order derivative terms in (3.9) vanish. Hence, it is natural to choose $\alpha_{0}^{\left(j\right)}=0$ and $\beta_{0}^{\left(j\right)}=0$ for any $j=\left\{-1,0,1\right\}$. At the upper boundary $x_{N_{x}}=x_{\max}$, we make the natural assumption that the solution $v$ behaves linearly in $x$, thus we choose $\beta_{N_{x}}^{\left(j\right)}=0$ for any $j=\left\{-1,0,1\right\}$, and we approximate the first-order derivative by the first-order backward finite difference scheme.
@@ 3.2.2 Summation term
    Here, $0_{P\times Q}$ denotes the null matrix of dimensions $P\times Q$, $\text{vec}\left(\cdot\right)$ denotes the vectorization of a matrix, $\sharp$ indicates the number of points of a given grid in one direction and $\Omega\in\mathbb{R}^{\sharp\mathbf{z}\times\sharp\mathbf{z}}$ is the matrix whose entries are the coefficients $\omega_{l_{1}l_{2}}$ defined by (3.5). For an example of a matrix $C$, we refer to Appendix C.
@@ 3.3 Temporal discretization: the semi-Lagrangian $\theta$-method
$Y^{n,0}=\begin{cases}V^{n-1}&\textrm{if }n=1,\\ 2V^{n-1}-V^{n-2}&\textrm{if }n=2,\\ 3V^{n-1}-3V^{n-2}+V^{n-3}&\textrm{if }n=3,\\ 4V^{n-1}-6V^{n-2}+4V^{n-3}-V^{n-4}&\textrm{if }n\geq 4.\end{cases}$ (3.30)
@@ 4 Numerical experiments
$\varphi\left(\xi\right)=\begin{cases}c\xi&\textrm{if }0\leq\xi\leq\xi_{\rm int},\\ x_{\rm int}+c\sinh\left(\xi-\xi_{\rm int}\right)&\textrm{if }\xi_{\rm int}<\xi\leq\xi_{\rm max},\end{cases}$
@@
where $K_{\nu}\left(\tau\right)=\frac{1}{2}\int_{0}^{\infty}y^{\nu-1}e^{-\frac{1}{2}\tau\left(y+y^{-1}\right)}dy$, for $\tau>0$, denotes the modified Bessel function of the second kind [^6] See Schoutens (2003, Appendix A)., $\left\langle x,y\right\rangle_{\rho}=x^{\top}\rho^{-1}y$ and $\left\|x\right\|_{\rho}=\sqrt{\left\langle x,x\right\rangle_{\rho}}$ is its induced norm. The constants $A_{\ell},B_{\ell}$, $C_{\ell}$ and $C^{\prime}_{\ell}$ in (2.6) are defined, with respect to $\left\|\cdot\right\|_{\rho}$, as
$A_{\ell}=2\alpha,\qquad B_{\ell}=\sqrt{\left\|\eta\right\|_{\rho}^{2}+2\lambda}-\left\|\eta\right\|_{\rho},$
$C_{\ell}\left(h\right)=\delta\frac{2^{\alpha}\Gamma\left(\alpha+1\right)}{\pi\sqrt{\det\left[\rho\right]}}e^{h\left\|\eta\right\|_{\rho}},$
$C_{\ell}^{\prime}\left(h\right)=C_{\ell}\left(h\right)\left\|\rho^{\frac{1}{2}}\right\|_{\rho}\left\|\rho^{-1}\right\|_{\rho}\left(h\left\|\eta\right\|_{\rho}+\left(A_{\ell}+2\right)\right).$
@@
    $z_{\max}^{\mathbf{II}}=h_{z}\cdot\text{ceil}\left[\frac{\sqrt{0.1}z_{\max}^{\mathbf{III}}}{h_{z}}-\frac{1}{2}\right]$, such that [^8] Whereas $z_{\max}^{\mathbf{II}}$ is not independent of $h_{z}$, the dependence is weak and the proof in Appendix B is readily adapted such that the convergence result of Proposition 3.1 remains valid. $z_{\max}^{\mathbf{II}}\rightarrow\sqrt{0.1}z_{\max}^{\mathbf{III}}$ as $h_{z}\rightarrow 0^{+}$. This means that the size of $R_{z}^{\mathbf{II}}$ is about 10% of the full integration domain $R_{z}$.
@@
    $z_{\max}^{\mathbf{III}}=\max\left\{\left\|z\right\|_{\infty}:z\in\mathbb{R}^{2},\ell\left(z\right)=10^{-8}\right\}$. Since the Lévy measure decays at least exponentially as $\left\|z\right\|\rightarrow\infty$, we ensure that $\ell\left(z\right)<10^{-8}$ for all $z\in\mathbb{R}^{2}$ such that $\left\|z\right\|_{\infty}>z_{\max}^{\mathbf{III}}$. Thus, we choose $z_{\max}^{\mathbf{III}}$ equal to 11.5010 for VG0, 2.1410 for VG1, 0.4172 for NIG0, and 0.8807 for NIG1.
@@
    $c$ is chosen such that $F=\max\left(65\%,\frac{x_{\rm int}}{x_{\rm max}}\right)$. In this way, approximately at least 65% of the spatial grid points in each given direction belong to the interval $\left[0,x_{\rm int}\right]$. Thus, we choose $c$ equal to 21.6164 for VG0, 67.1487 for VG1, 55.0673 for NIG0, and 45.0189 for NIG1.
@@
    $N_{y}^{-}=\text{ceil}\left[-\frac{1}{h_{z}}\ln\left(x_{1}\right)\right]+N_{y}^{*}$ and $N_{y}^{+}=\text{ceil}\left[\frac{1}{h_{z}}\ln\left(x_{\rm max}\right)\right]+N_{y}^{*}$ for some given $N_{y}^{*}\in\mathbb{N}$. This choice minimizes the need for extrapolation in (3.18) as it is necessary to extrapolate just to the grid points $x_{m_{1}m_{2}}$ with either $m_{1}=0$ or $m_{2}=0$. This is done in a linear fashion. In (3.17), as well as (3.13), we set $v\left(x,t\right)=0$ whenever $x\notin R_{x}$.
@@
    $N_{y}^{*}$ is taken as the minimal $n\in\mathbb{N}$ such that the maximal prime factor of $\sharp\textrm{in}=N_{y}^{-}+N_{y}^{+}+2N_{z}$ is at most 7. This is beneficial for the efficiency of the FFT.
@@
    The tolerances for the fixed-point iteration and BiCGSTAB are set to $10^{-7}$ and $10^{-14}$, respectively.
@@
$\left(90,90\right)$ 12.6540 10.1079 11.4067 11.5833
$\left(90,100\right)$ 10.6127 5.8453 7.8724 8.1532
$\left(90,110\right)$ 9.0142 3.0171 5.1023 5.4661
$\left(100,90\right)$ 10.4066 5.7627 7.8897 8.0913
$\left(100,100\right)$ 8.8020 2.9029 5.1186 5.3956
$\left(100,110\right)$ 7.5314 1.3889 3.1156 3.4314
$\left(110,90\right)$ 8.6186 2.8062 5.1393 5.3384
$\left(110,100\right)$ 7.3468 1.3058 3.1326 3.3739
$\left(110,110\right)$ 6.3294 0.6012 1.7937 2.0401
provides the numerical option prices for various points $x$ around $\left(K,K\right)$ and all parameter sets from Table 1. Here $N_{x}=800$ for VG0, VG1, and NIG1, while for NIG0 we used $N_{x}=400$.
We next investigate the convergence behaviour of the numerical scheme. Let $\mathbf{x}_{N}$ denote the set of spatial grid points if $N_{x}=N$. For $x\in R_{x}$, let $\tilde{u}(x;N)$ denote the approximation of the exact solution value $u(x,T)$ obtained by the numerical scheme if $N_{x}=N$. More precisely, the vector $V^{N_{t}}$ generated by (3.28)-(3.29) yields the approximation on the spatial grid $\mathbf{x}_{N}$ and cubic interpolation is employed whenever $x\notin\mathbf{x}_{N}$. We consider $\tilde{u}(x;N)$ with $N=400$ as the reference solution and study for $50\leq N\leq 200$ the total error defined by
@@
displays the total errors for all four parameter sets from Table 1. The quantity $p$ in the legend denotes the numerical order of convergence, which is computed by linear regression. Clearly, the favourable result is found that the numerical scheme achieves second-order convergence for each set of parameters. This is in agreement with the theoretical order of convergence for the discretization of the integral term given by Proposition 3.1 in the case of the two VG sets (where $A_{\ell}=0$). On the other hand, in the case of the two NIG sets (where $A_{\ell}=1$), the theoretical order of convergence is only one. At present we have no clear explanation why, for the two NIG sets, second-order convergence is observed in the numerical experiments. We shall leave this interesting question for future research.
@@ 5 Conclusions
In this paper, we have developed an effective numerical method for the valuation of European options under two-asset exponential Lévy models with particular attention to the infinite-activity case.
Our method is based upon the ideas in Wang et al. (2007) for the one-asset case. A key part of our method is the tailored discretization of the non-local double integral term, designed to handle singular measures under mild assumptions. The discretized integral term can subsequently be efficiently evaluated by FFT. For the discretization in time, the semi-Lagrangian Crank–Nicolson method is employed with a fixed-point iteration on the integral part.
Numerical experiments for put-on-the-average options under Normal Tempered Stable processes indicate that our method achieves favourable second-order convergence whenever the exponential Lévy model has finite-variation. A relevant rigorous theoretical result is proved on the convergence behaviour of the discretization of the integral term.
The numerical method derived in this paper is easily adapted to many other European options on two assets, such as spread options in energy markets.
@@ Appendix A $d$-dimensional Normal Tempered Stable process
In our context, we will use this process to define the jump component of the logarithmic return in asset prices, i.e., we choose
@@ A.1 Tempered Stable subordinator
Lévy measure $\ell_{G}\left(z\right)=\mathbb{I}_{z>0}\delta e^{-\lambda z}z^{-1-\alpha}$
Characteristic exponent $\psi_{G}\left(z\right)=\begin{cases}-\delta\ln\left(1-iz\lambda^{-1}\right)&\text{if }\alpha=0\\ \delta\Gamma\left(-\alpha\right)\left(\left(\lambda-iz\right)^{\alpha}-\lambda^{\alpha}\right)&\text{if }\alpha\in\left(0,1\right)\end{cases}$
@@
Density function $f_{G}\left(z\right)=\begin{cases}\mathbb{I}_{z>0}\frac{\lambda^{\delta}}{\Gamma\left(\delta\right)}z^{\delta-1}e^{-\lambda z}&\text{if }\alpha=0\\ \mathbb{I}_{z>0}\delta z^{-\frac{3}{2}}e^{-\left(\sqrt{\lambda}z-\delta\sqrt{\pi}\right)^{2}z^{-1}}&\text{if }\alpha=\frac{1}{2}\\ \text{not known analytically}&\text{else}\end{cases}$
@@ A.2 Normal Tempered Stable process
Consider $B\left(t\right)=\eta t+\sqrt{\rho}\,W\left(t\right)$, where $W$ is a standard $d$-dimensional Wiener process, $\eta\in\mathbb{R}^{d}$ and $\sqrt{\rho}$ is obtained by the Cholesky decomposition of a given positive definite symmetric matrix $\rho$, i.e. $\rho=\sqrt{\rho}\cdot\sqrt{\rho}^{\top}$. Adapting the results presented in Barndorff-Nielsen et al. (2001) and Rocha-Arteaga & Sato (2019, Chapter 4) [^9] The authors consider the more general case where the characteristic exponent of $L$ is defined as $\psi_{L}\left(\tau\right)=\int_{\mathbb{R}^{d}}\left(e^{i\tau^{\top}z}-1-i\tau^{\top}z\mathbb{I}_{\left\|z\right\|<1}\right)\ell_{L}\left(dz\right)$, while we consider the case where $\psi_{L}\left(\tau\right)=\int_{\mathbb{R}^{d}}\left(e^{i\tau^{\top}z}-1-i\tau^{\top}z\right)\ell_{L}\left(dz\right)$., we define a $d$-dimensional Normal Tempered Stable process as
@@
Lévy measure $\ell_{L}\left(z\right)=\delta\Phi\left(z\mid\alpha,0\right)$
Characteristic exponent $\psi_{L}\left(z\right)=\begin{cases}-\delta\ln\left(\frac{\lambda-iz^{\top}\eta+\frac{1}{2}z^{\top}\rho z}{\lambda}\right)-iz^{\top}c&\text{if }\alpha=0\\ \delta\Gamma\left(-\alpha\right)\left(\left(\lambda-iz^{\top}\eta+\frac{1}{2}z^{\top}\rho z\right)^{\alpha}-\lambda^{\alpha}\right)-iz^{\top}c&\text{if }\alpha\in\left(0,1\right)\end{cases}$
@@
Density function $f_{L}\left(z\right)=\begin{cases}\frac{\lambda^{\delta}}{\Gamma\left(\delta\right)}\Phi\left(z+c\mid-\delta,0\right)&\text{if }\alpha=0\\ \delta e^{2\delta\sqrt{\lambda\pi}}\Phi\left(z+c\mid\frac{1}{2},\delta^{2}\pi\right)&\text{if }\alpha=\frac{1}{2}\\ \text{not known analytically}&\text{else}\end{cases}$
@@
$\Phi\left(z\mid a,b\right)=2\sqrt{\frac{\left(\left\|\eta\right\|_{\rho}^{2}+2\lambda\right)^{a+\frac{d}{2}}}{\left(2\pi\right)^{d}\det\left[\rho\right]}}\frac{K_{a+\frac{d}{2}}\left(\sqrt{\left(\left\|\eta\right\|_{\rho}^{2}+2\lambda\right)\left(\left\|z\right\|_{\rho}^{2}+2b\right)}\right)}{\left(\sqrt{\left\|z\right\|_{\rho}^{2}+2b}\right)^{a+\frac{d}{2}}}e^{\left\langle\eta,z\right\rangle_{\rho}}$ (A.2)
where $K_{\nu}\left(\tau\right)=\frac{1}{2}\int_{0}^{\infty}y^{\nu-1}e^{-\frac{1}{2}\tau\left(y+y^{-1}\right)}dy$, for $\tau>0$, denotes the modified Bessel function of the second kind (see Schoutens (2003, Appendix A)), $\left\langle x,y\right\rangle_{\rho}=x^{\top}\rho^{-1}y$ and $\left\|x\right\|_{\rho}=\sqrt{\left\langle x,x\right\rangle_{\rho}}$ is its induced norm. We conclude this appendix with the following proposition.
@@
Consider a Lévy measure $\ell$ over $\mathbb{R}_{*}^{d}=\mathbb{R}^{d}\setminus\left\{0\right\}$. Assume that there exist constants $A_{\ell}$ and $B_{\ell}$, and for any given $h>0$ two constants $C_{\ell}\left(h\right)$ and $C_{\ell}^{\prime}\left(h\right)$ such that
[…]
where $\ell_{z}^{\left(j\right)}$ denotes the $j$-th partial derivative of $\ell$ with respect to $z$, with $j=1,2,\ldots,d$.
@@
$A_{\ell}=2\alpha,\qquad B_{\ell}=\sqrt{\left\|\eta\right\|_{\rho}^{2}+2\lambda}-\left\|\eta\right\|_{\rho},$
$C_{\ell}\left(h\right)=\delta\frac{2^{\alpha+\frac{d}{2}}\Gamma\left(\alpha+\frac{d}{2}\right)}{\sqrt{\left(2\pi\right)^{d}\det\left[\rho\right]}}e^{h\left\|\eta\right\|_{\rho}}\,$,
$C_{\ell}^{\prime}\left(h\right)=C_{\ell}\left(h\right)\left\|\rho^{\frac{1}{2}}\right\|_{\rho}\left\|\rho^{-1}\right\|_{\rho}\left(h\left\|\eta\right\|_{\rho}+\left(A_{\ell}+d\right)\right).$